OpenStudy (friendzone):

There are 15 cards, numbered 1 through 15. If you pick a card, what is the probability that you choose an odd number or a two? A) 2 3 B) 3 5 Eliminate C) 8 15 D) 8 225

OpenStudy (anonymous):

3/5 .... P(odd)=8/15 P(2)=1/15 P(odd)+p(2)=8/15+1/15=9/15=3/5